OpenStudy (babyslapmafro):

Can someone please show me where I went wrong in the following problem... Use an appropriate local linear approximation to estimate the value of the given quantity. cos(31*)

OpenStudy (babyslapmafro):

Approximation: cos(pi/6) x=(pi/6) f(x)=cos(pi/6) f'(x)=-sin(pi/6)=-1/2

OpenStudy (babyslapmafro):

point: (pi/6,cos(pi/6)) y-cos(pi/6)=-1/2(x-pi/6)

OpenStudy (babyslapmafro):

y=-1/2(cos(31pi/180)-pi/6)+cos(pi/6)=.69924

OpenStudy (babyslapmafro):

The answer in the book is 0.8573...where did I go wrong?

OpenStudy (tkhunny):

Well, let's be a little more deliberate about it. y = cos(x) sqrt(3)/2 = cos(30º) dy = -sin(x)*dx dy = -sin(30º)*(pi/180) = (1/2)*(pi/180) y + dy = cos(x) - sin(x)dx y + dy = sqrt(3)/2 - (1/2)*(pi/180) = 0.8573